Kindly answer these problems please because i am still quite confused in answering these problem using system of linear equations.
3) Working together, James and Chris painted their gate in 3 hrs. and 20 min. Working alone, Chris woul need 5 hrs. longer than James to finish the job. How long would it take each of them alone to do the job?
Thank you in advance! Hoping for your immediate reply.

Answer by mananth(16946)

James takes x hours
1/x of the job he does in 1 hour
chris takes x+5 hours
1/(x+5) of the job he does in 1 hour.
...
Together they take 3 hours 30 minutes = 3 1/3 hour = 10/3 hours
..
so together they do 3/10 of the job in 1 hour
...
1/x +1/(x+5) = 3/10
multiply by LCD 10x(x+5)
10x+50+10x = 3x(x+5)
20x+50 = 3x^2+15x
3x^2+15x-20x-50=0
3x^2-5x-50=0
3x^2-15x+10x-50=0
3x(x-5)+10(x-5)=0
(x-5)(3x+10)=0
x= 5 hours --- james
Chris ---- 10 hours